OpenStudy (anonymous):

what is 8.05 as a fraction and percent?

OpenStudy (ahvei):

8 5/100 is the fraction

OpenStudy (anonymous):

i think its 8 1/20 and 80.5% is that right ?

OpenStudy (anonymous):

Hi crazy_girl, You can take 8.05 and multiply by 100/100 (that's just multiplying by 1, so it doesn't change the number --- just the "name" of the number!) Then, 8.05 * 100/100 = 805/100; it can be reduced, if desired. This name makes it easy to go to the percent name, since 1% = 1/100: 805/100 = 805*(1/100) = 805% So --- the number 8.05 has *lots* of different names; and different names are good for different things you might want to do with the number!

OpenStudy (anonymous):

so did o get it right

OpenStudy (anonymous):

And... one of my web pages might be helpful for you; it talks about changing decimals to percents: http://www.onemathematicalcat.org/algebra_book/online_problems/dec_to_percent.htm It talks about a *really easy way* to go from a decimal to a percent. There's lot of practice at the bottom, if you want feedback about if you've got the idea down... hope it helps!

OpenStudy (anonymous):

ok